### 07 Dec 2023 v1

# Cryogenic platform to investigate strong microwave cavity-spin coupling in correlated magnetic materials

Abstract
We present an experimental demonstration of loop-gap resonators (LGRs) for continuous-wave electron spin resonance (ESR) spectroscopy at cryogenic temperatures and to investigate the hybridization of solid-state magnetic materials with microwave cavity modes. The simple experimental setup, implemented in a conventional cryogenic platform for materials research, allows to obtain ESR spectra to temperatures as low as K. The versatility of this approach is demonstrated through experiments on two transition-metal based Heisenberg antiferromagnets, CuSO5HO and MgCrO, showcasing simple measurements of their gyromagnetic factor and magnetic susceptibilities. Our work also demonstrates the realization of strong microwave cavity-spin coupling in CuSO5HO and delves into the challenges of fitting ESR spectra under strong hybridization conditions, underscoring the significance of proper calibration and stabilization.
Keywords: electron spin resonance spectroscopy, quantum magnetism, spin liquids, spin-polariton, microwave cavity
1 Introduction
Electron spin resonance (ESR) spectroscopy is a versatile experimental technique to study and control the spin dynamics of dense and dilute material systems through microwave-induced spin transitions under an applied magnetic field. The technique can be implemented in continuous wave (CW) and pulsed forms, providing direct determinations of a system’s -tensor, relaxation times, and magnetic susceptibilities. The low-frequency nature of ESR in the X-band (), requires microwave resonators. This can be seen through the relative frequency () of the induced transitions compared to the thermal energy of the samples at equilibrium (). In the regime, thermal excitations lead to weak population differences and absorption between magnetic states [1]. By enhancing light-matter interactions, resonators produce a corresponding increase in microwave sensitivity, allowing small changes in sample inductance to be detected.
Recent experimental work has focused on studying magnetic quantum phases and hybrid quantum states that emerge from spin-microwave photon interactions in cavities. Applications range from quantum information and material control [2, 3] to understanding exotic magnetism [4, 5]. But investigating the low-energy spectral response of correlated materials with electron-spin-based resonance techniques has long offered unique advantages, with experiments testing exquisite theories of magnon fractionalization into spinon pairs in low-dimensional spin systems [6, 7] and spin-orbital quantum magnets [8, 9]. Sensitivity to the spinon continuum is associated with an increasing resonance linewidth with decreasing temperature. However, in practice, thermal fluctuations obscure the observation of these collective quantum phenomena. Gaining a more precise picture thus requires performing measurements at sub-Kelvin temperatures beyond the limit of most commercial and very high magnetic-field experimental set-ups.
Additionally, strong spin-lattice coupled materials typically show no visible ESR spectra due to their rapid relaxation rates. In these cases, phonons dissipate absorbed microwave energy, resulting in broad spectral peaks that only become observable in the low-temperature, single-Kelvin regime, where relaxation rates slow and lattice vibrations die out [3]. This is the case with many rare-earth-based quantum magnets.
This manuscript reports a straightforward cryogenic implementation of loop-gap resonators (LGRs) to investigate the feasibility of different magnetic quantum materials to hybridize with a microwave polariton mode down to K. Our work integrates past results demonstrating the utility of LGRs to control hybrid quantum information science systems [2] and to obtain traditional ESR spectra on molecular magnets at single-Kelvin temperatures using a commercial Quantum Design Physical Property Measurement System (PPMS) [10]. The flexibility of LGRs and our low-cost demonstration experimental setup not only enables C-band ESR spectroscopy (– GHz) inside a cryostat environment but also serves as a platform to characterize the potential for strong spin-polariton coupling between microwave resonators and spin-dense, correlated materials. We demonstrate our approach on two magnetic insulators described by a Heisenberg exchange Hamiltonian and hosting quantum and classical spin-liquid correlations, respectively. We also review the relevant technical concepts in the hope of helping future successful implementations by other laboratories and paving the way for our future implementation in a dilution refrigerator reaching the sub-Kelvin regime.
2 Design Experimental Set-Up and Method
2.1 Resonators
A critical component in conventional ESR spectroscopy and many hybrid quantum systems setups is the microwave resonant structure that hosts the sample or system of interest. Maximizing microwave sensitivity in ESR typically requires maximizing the product of the resonator quality factor () and filling factor (). For this study, we utilized lumped-element resonators known as loop-gap resonators (LGRs). These structures were first developed in the 1980’s for nuclear [13] and electron [14] magnetic resonance studies in sub X-band frequencies. While LGRs are limited in their accessible frequency range to to GHz for the geometries considered here, and like many other lumped element resonators, operating with much smaller quality factors (), their significant filling factors () and homogeneous field (), makes them ideal for strong coupling to bulk magnetic materials in low-temperature environments.
The specific design we adopted uses a three-loop, two-gap layout structure first described in 1984 for ESR and electron-nuclear double resonance spectroscopy (ENDOR) spectroscopy [11], and recently utilized in experiments investigating quantum transduction phenomena [2, 3]. One of the primary benefits of a multi-loop and gap resonator is the ability to manage and contain the returning magnetic flux pathways in a microwave experiment, making them ideal to use in cryostats and dilution refrigerators where large superconducting magnets limit the free space available to construct an experiment.
The geometric parameters that determine an empty LGR resonance frequency and quality factor are shown in Fig. 1(a) following their previous description in Ref. [11]. For the experiments carried out in this work, aluminum LGRs were created with a manual milling machine in two halves to create the sized gaps. The two halves were then attached with a silver-based conductive epoxy to form the full resonator. While this manufacturing approach does yield lower quality factors than precision-machined LGRs [2], we find them adequate to observe the necessary spectra for our demonstration experiments; future studies will be carried out on higher-quality resonators.
While modeling a desired resonator sample size and frequency is possible with the design equations in the early literature [11], this only resulted in predictions accurate within 1 GHz of the experimental resonance frequency. Instead, we utilized COMSOL Multiphysics to simulate and find the eigen-frequencies and modes for many of our resonator designs. In the following, we refer to as the main resonance frequency of interest for our experiments.
2.2 Sample Selection
To explore strong cavity-material coupling in samples that are known to host correlated spin physics, we chose to investigate two transition-metal based Heisenberg antiferromagnets: the one-dimensional quantum liquid CuSOHO and the three-dimensional classical spin-liquid MgCrO. While these materials have been well studied using a host of experimental techniques, from traditional and high-field ESR to inelastic neutron scattering experiments [15, 16], our experiments investigate for the first time the ability of these systems to strongly hybridize with a cavity mode over a range of temperatures. These samples were chosen due to their well-characterized ESR response, availability as high-quality large single-crystals fitting our LGR, Fig. 1(b), and distinctly different magnetic behavior regarding magnetic excitation bandwidth and the role of quantum correlations.
2.3 Microwave Generation, Detection and Transport in Cryogenic Environment
A R60 1-Port 6 GHz Vector Network Analyzer (VNA) from Copper Mountain Technologies was used as the microwave source and detector. A diagram of the finalized microwave setup used is shown in Fig. 1(d). The signal from the VNA was fed into a stub tuner using a procedure detailed in Ref. [17]. A semi-rigid coax cable was then used to transmit the microwave signal through a hermetically sealed union to the bottom of a Quantum Design cryogen-free Physical Property Measurement System (PPMS) equipped with a 14 T superconducting magnet and reaching base temperature of K. The experiment itself used Quantum Design’s Multi-Function Probe (MFP), which provided enough space for both the sample plus LGR assembly, the coupling structure and the right-angle mount that secured the LGR with the loops perpendicular to the applied DC field.
2.4 Microwave Coupling
Mutual inductive coupling was utilized between the LGR body and the coaxial transmission line, see e.g. Fig. 1(c). In practice, two different coupling setups were used, labeled setup A and B in Fig. 1(e) and (f), respectively. Maximizing the microwave power transferred to the resonator is important and characterized by the degree to which critical coupling is achieved. Critical coupling, or matching, occurs when the impedance of the transmission line equals that of the resonator. This condition minimizes reflected microwave power, and various methods to create these matching networks for LGRs are well documented [18]. The coupling for each setup was optimized by analyzing the Smith chart for the VNA trace section of interest. In setup A, we attempted to critically couple by physically moving the inductive loop on the end of the coax line on the bench top. While this allowed us to get a detectable signal, it was distinctively over-coupled and limited our ability to fine-tune the setup inside the cryostat. We improved flexibility using a stub tuner, which controlled the impedance of the coax line to match the resonator at its resonance frequency, as seen in figure 1(d). This facilitated in-situ tuning before the ESR field sweeps inside the cryostat. In setup B, a printed circuit board (PCB) coupling loop was used with an additional FR4 material as a spacer between the LGR body and the copper trace. This provided a more stable configuration that experienced minimal drift as the experiment was cooled from room temperature inside the cryostat. More details of our set-up are provided in the caption of Fig. 1.
2.5 Sample Preparation and Orientation
For high sensitivity of ESR response, including at room temperature, the resonators and the samples had to match in size to ensure a high filling factor. For CuSOHO, this entailed polishing a large crystal down to the size of the LGR sample loop, while floating-zone grown MgCrO crystals [19] were observed to already fit well within the first LGR sample radius of 2.3 mm, Fig. 1(b). While CuSOHO fit well within the resonator, the exact orientation of the crystal was not determined, as an exact -factor determination was second to observing an initial ESR response for the experimental setup. The sample was adhered to the resonator body with Kapton tape, which crossed over one side of the sample loop, and can be seen behind the CuSOHO crystal in Fig. 1(b). For MgCrO, the nearly Heisenberg nature of the magnetism meant that precise sample alignment was also not necessary for this demonstration study.
2.6 Sample-Resonator Mounting
The sample-resonator union was mounted onto a 3D-printed holder with an open rectangular slot to firmly secure the bottom section of the resonator away from the sample loop and return flux paths. This slot was then mounted onto an electric transport puck built for the PPMS multi-function probe (MFP) so that the entire setup was attached to the rest of the probe. The bottom of the 3D printed slot was attached with super glue to ensure the setup would not move during the experiment. The next step was to feed through a semi-rigid coax cable, threading it through the center hole on the head section of the MFP. One of the more tedious parts of the design was attaching the coupling loop to the bottom of this coax line. For the room temperature setup described here (setup A), the coupling loop was made of a stripped piece of coax where the central conductor was twisted and soldered to the outer conductor to form an inductive coupling loop. This coax termination transmitted the VNA signal into the loop-gap resonator, which then interacted with the sample. A combination of Teflon and adhesive was used to secure the coupling loop to the side of the MFP probe sample space, whose orientation is shown in Fig. 1(e). Stabilizing the coupling loop was a crucial part of taking these first room temperature measurements, as any movement would produce a background signal superimposed on top of the material spectra. It should be noted that while the varnish method of stabilizing the resonators and coupling was useful, future design will utilize a more standard designed structure with brass screws and mounting components.
2.7 Measurement Protocol
The DC magnetic field of the PPMS was applied along the vertical direction and was then swept between Oe to Oe in steps of 10 Oe. In addition, the probe was swept between Oe and Oe to reveal any field asymmetries in the setup’s response. At each point, the field was allowed to equalize for a few seconds before a trace was taken, where is akin to the power reflection coefficient at a given frequency. The raw data consisted of csv files of the VNA traces taken over a 1.5 GHz span centered on each sample-cavity resonance frequency. To get an overview of the data and observe if hybridization of the cavity polaritons and material spins occurred, the ESR spectra would be visualized as a 3D colormap by plotting the vs. frequency (GHz) traces as a function of . This technique follows those used by other studies [2, 3, 20].
3 Results
3.1 Room Temperature
The room-temperature results (“diagnostic plots” using setup A) for both CuSOHO and MgCrO in their uncorrelated paramagnetic regime are presented in Fig. 2(a) and Fig. 2(b), respectively. Both compounds show a clear ESR response at room temperature, as visible, for CuSOHO, from the broadening by around MHz of the VNA frequency lineshape beyond the baseline full-width at half-maximum of MHz and slight shift of the peak frequency with respect to . For MgCrO, broadening by around MHz occurred from the MHz baseline. Calculating the -factor and plotting versus the applied magnetic field yields an unaligned -factor of for CuSOHO, while the line-width of the ESR response was measured to be Oe full-width-half maximum, centered at a resonance field of Oe. MgCrO, on the other hand, displayed a much broader response, characteristic of its highly correlated paramagnetic spins, with a Weiss constant of K for MgCrO [16] compared to K for CuSOHO. The measured -factor for MgCrOis , while the line-width was Oe FWHM, centered at Oe. Both -factor measurements agree with previously reported results [21, 12]. The cavity’s quality factor (measured off-resonance) was approximately 320 in these conditions. While this low-quality factor can hinder sensitivity, the large filling factor () can compensate for this deficiency, allowing us to reach the regime of strong spin-photon hybridization at low temperatures.
3.2 Data Analysis
Results obtained at low temperatures using setup B are presented for CuSOHO and MgCrO in Fig. 3 and 4, respectively. Before discussing these results, we present our data analysis approach in more detail, given that several strategies to reduce the raw data from the VNA are possible once the cavity and sample excitation mode hybridize.
One of the key advantages of the VNA method of ESR detection is being able to separate both the real () and imaginary () components of complex susceptibility simultaneously [20, 22]. In conventional ESR, only the reactive () or dispersive () component is measured on a single resonance frequency. This can be directly detected via the reflected microwave power, or the signal can commonly be converted to a DC voltage via a Schottky diode, and is related to spectrometer parameters of operation through [23]
where is the characteristic impedance of the transmission line (typically ), the input microwave power to the sample-resonator system, and is the loaded quality factor defined as with the sample inside the resonator. Many commercial EPR spectrometers utilize an automated frequency control (AFC) circuit to stabilize the cavity resonance and optimize the reflected microwave output for a specific sample size [24]. It is possible to approximate this frequency stabilized resonance and thus observe the data in a traditional ESR format by tracking how changes with DC magnetic field, similar to the techniques of other studies [25, 10]. There exist many analysis techniques originating from the superconducting circuits community to accurately determine the quality factor of a resonating circuit from a VNA trace, whether it be a Smith or plot, and reduce the effects of background signal from coax cables and environmental factors [26].
3.3 Cryogenic Response and Hybridization
It becomes difficult to accurately fit a VNA trace when the cavity-spin system becomes hybridized. This situation occurs in setup B when CuSOHO is cooled below K, see Fig. 3(a) which displays a strong avoided crossing phenomena between the cavity mode and the Zeeman split paramagnetic excitation of the sample reaching up to Oe. In that situation, the original central resonance frequency disappears, and two weaker hybridized modes emerge, which become difficult to fit and determine a full width at half max (FWHM) value to extract the -factor. Compounding this effect is the periodic background signal coming from the length of coax cable used in the experiment. For this reason, the -factors for each trace, shown in Fig. 3(b) are determined from a simplified fit that extracts the maximum value of at the resonance position and collects the nearest data points at half that value to determine the frequency FWHM and the -factor. This procedure neglects the raised baseline produced by the background coax-signal, resulting in a broader FWHM determination than the actual value. However, the negative impact is only observed at the resonance matching condition when the signal broadens to its most considerable extent for CuSOHO and results in the -factor appearing to reach a zero value. This quick drop to at resonance matching can be seen for the three lowest temperatures in Fig. 3(b). While plotting is primarily used in traditional ESR, here it allows a depiction of how spin-polarition coupling strength is increasing, as our zeroth-order Lorentzian fitting program outlined in the text can accurately obtain the VNA dip linewidth up to the regime where hybridization makes fitting with the coax background impossible and (as seen in the plot).
In contrast to CuSOHO temperature dependence results for MgCrO Fig 4(a), display no strong hybridization. The -factor vs. magnetic field plots shown in Fig 4(b) allow for a traditional ESR interpretation, and the decreasing paramagnetic response as the sample approaches is consistent with previous reports for MgCrO[21].
Usually, matching at the resonance condition between the DC field and the microwave frequency produces a decreased loaded quality factor that can be measured as the ESR response. This behavior was observed in our set-up A, see Fig. 2. This result is usually explained in textbooks [24] (and observed in experiments [25]) as more energy is dissipated by flipping paramagnetic spins or generating magnons, resulting in a quality factor decrease. However, this is not always the case, as can be seen in Fig. 4(b), where the quality factor increases upon reaching the resonance condition. Instead, this data can be better understood as the magnetic material changing the reactance and thus the coupling of the cavity-sample system to the transmission line. Whether or not the magnetic material causes the quality factor to increase or decrease depends on how close the off-resonance Lorentzian is to critical coupling and if the resonance frequency change associated with matching takes the peak closer or further to the critical coupling condition. If the system is critically coupled or close to it, reaching resonance will always decrease the quality factor. Still, an increasing quality factor can be observed when the setup is far enough from critical coupling. This shows why it is essential to properly calibrate and stabilize the microwave setup. The magnitude of the ESR response is still sample-dependent, so even off of the critical coupling condition, we are able to obtain significant results. Once the stub tuner was adjusted to reach closer to critical coupling, the same temperature-dependent results in Fig. 4(b) were repeated, but this time with a quality factor that decreased as expected (results not show).
In Fig. 3(c) shows every VNA trace taken between 1500 to 1800 Oe for CuSOHO, where strong hybridization is seen to occur between the two grey arrows located in the bottom plot of Fig. 3(a).
These traces have each been modified by subtracting the same background trace from each profile. This is to remove the repeated coax signal previously mentioned and observe the pure cavity-sample response of each. However, it can be seen that periodic irregularities still exist in that there are slight protrusions above 0 dB at the frequencies of 4.950 GHz, 5.025 GHz, and 5.100 GHz. It should be noted that it was confirmed that this background signal was from the coax cable by using the well-known expression:
which produces a length of 1.4 m for Teflon’s dielectric constant, and a frequency dip separation of 75 MHz. This matched the length of the cable used to within a centimeter, confirming the signal’s origin. The background trace was created by stitching together the trace at 1600 Oe from the lowest frequency collected to the central resonant frequency ( ) with the trace at 1700 Oe from the central resonant frequency to the highest frequency collected. This allows the background signal to be reconstructed by adding opposing trace sections near the avoided level crossing that reveal opposite sections of the true background that the sample-LGR resonance frequency is superimposed onto. For VNA traces of MgCrO over its entire swept spectra in 100 Oe intervals, shown in Fig. 4(c), no background trace is subtracted since no strong hybridization occurs, meaning the frequency region of interest does not span a large enough region for the coax signal to obscure any results. With the background trace subtracted for CuSOHO, the faint hybridized modes are seen for traces around the center of Oe.
3.4 Interpretation
Overall, the primary difference between the two sample spectra studied down to K is that CuSOHO is seen to strongly hybridize with the cavity mode, increasing the degree of hybridization with a decreasing temperature, while MgCrO shows no hybridization with the resonator environment despite experimental considerations such as filling factor, resonance frequency, and mode-volume confinement being the same.
The coupling strength between paramagnetic spins and the microwave cavity is provided as [3]:
where is the number of magnetic ions in the sample and is the cavity mode volume. For CuSOHO at K, the coupling strength is estimated to be on the order of MHz, but a more precise estimate would require modeling the spin-cavity Hamiltonian. The lack of hybridization for MgCrO is speculated to arise from its more strongly correlated paramagnetic spins that produce a broad ESR response even at room temperature, however, the mere observation of sample-cavity coupling differences between these two samples demonstrates that our setup probes both the hybridization and magnetic response across a wide range of temperatures in a relatively simple experimental architecture.
4 Conclusion
This manuscript reports the design and implementation of loop-gap resonators (LGRs) for electron spin resonance (ESR) on quantum materials and hybrid quantum system studies of hybrid microwave polariton to quantum spin hybridization. The experimental setup is low-cost and implemented in cryogenic equipment commonly found in laboratories worldwide.
First experiments were conducted on single-crystals of paramagnetic salts down to K. Although the precise interpretation in terms of microscopic models is still under development, our findings indicate that sample-embedded loop gap resonators (LGRs) with high filling factors serve as unique and flexible probes for studying magnetic susceptibilities and strong spin-polariton coupling in hybrid quantum systems. Expanding upon previous studies that validate the viability of LGRs [2, 3, 27], our work applies them to investigate spin-liquid candidate materials of interest in quantum materials research. We highlight that this technique bridges the gap between well-established, traditional magnetic resonance studies and novel investigations of hybrid-quantum systems and heterostructures at low temperatures.
Our report hopes to assist other experimental groups in developing effective microwave resonance methods for studying bulk magnetic samples. Our method not only facilitates the acquisition of magnetic susceptibility data but also enables the accessible modification of the spin’s electromagnetic environment by inducing strong spin-polariton modes. By implementing our two setups on a widespread controllable platform, the Physical Property Measurement System (PPMS), along with accessible microwave equipment (stub tuner and network analyzer), and by providing a detailed step-by-step analysis process, including aspects not commonly reported in the literature, we contribute to the broader understanding of these techniques.
In our future research endeavors, we plan to enhance our experimental setup by introducing precision-machined resonators, optimizing the microwave bridge, and conducting experiments at lower temperatures in a dilution refrigerator. This upgraded configuration goes beyond mere integration with a network analyzer; it promises heightened control and versatility. Additionally, our upcoming studies aim to explore the time-domain response of the strongly coupled states identified in this work. We expect that these advancements will not only contribute to our understanding of an effective and accessible technique but also drive progress in the fields of condensed matter and quantum information.